# The torus plateau for the high-dimensional Ising model

### v1

@@ 3.2 A percolation exploration
The construction of $\mathcal{C}(0)$ is as follows. We fix an ordering of the vertices in $\mathcal{G}$.
@@ 3. Step 2
3.3 A coupled cluster exploration
We now apply the exploration introduced in Section 3.2 to a well-chosen percolation configuration induced by random currents. This allows us to perform switching between a current on a large finite subgraph $\mathcal{G}\subset\mathbb{Z}^{d}$ and a current on $\mathbb{T}_{r}$.
Let $\mathbf{m}$ be a current on $\mathcal{G}$ and $\mathbf{n}$ be a current on ${\mathbb{T}_{r}}$. We write $\hat{\mathbf{n}}$ for the lift of $\mathbf{n}$ to $\mathcal{G}$, defined by
$\hat{\mathbf{n}}(e)=\mathbf{n}(\pi e)\qquad(e\in\mathcal{G}),$ (3.10)
where $\pi:{\mathbb{Z}^{d}}\to{\mathbb{T}_{r}}$ is the canonical projection map. The current $\mathbf{m}+\hat{\mathbf{n}}$ induces a percolation configuration $(\mathds{1}_{(\mathbf{m}+\hat{\mathbf{n}})_{e}>0})_{e\in E}$ on $\mathcal{G}$. We run the exploration of Section 3.2 on this configuration and produce $(\mathsf{B},\mathsf{G},\mathfrak{B},\mathfrak{W})$ and $\mathcal{C}(0)=(\mathsf{B},\mathfrak{B})$.
@@ 3.3 A coupled cluster exploration
Let $d\geq 1$, $r\geq 3$, $x\in\mathbb{T}_{r}$, and fix finite $\mathcal{G}\supset\Lambda_{2r^{d}}$. Let $\mathbf{m}\in\Omega_{\mathcal{G}}$, and let $\mathbf{n}\in\Omega_{\mathbb{T}_{r}}$ with $\partial\mathbf{n}=\{0,x\}$. Then the exploration $(\mathsf{B},\mathsf{G},\mathfrak{B},\mathfrak{W})$ of $(\mathds{1}_{(\mathbf{m}+\hat{\mathbf{n}})_{e}>0})_{e\in E}$ satisfies the following properties:
@@
    $\hat{\mathbf{n}}_{e}=0$ for every representative of $e\in\mathfrak{W}$.
@@
We use the exploration to switch sources between currents on $\mathcal{G}$ and ${\mathbb{T}_{r}}$.
@@
In particular, for $x\in{\mathbb{T}_{r}}$, $A=\{0,x\}$, and $z=x^{\prime}$ with $\pi z=x$,
@@
By conditioning on $(\mathfrak{B},\mathfrak{W})$, we have
@@
where $B\ni z$ means that $z$ is incident to some edge in $B$. Given an admissible pair $(B,W)$ (i.e., realisable by the exploration and such that $B\ni z$), we make the decomposition
@@
$\mathbf{n}$ $=\mathbf{n}_{1}+\mathbf{n}_{2},\qquad$ $~{}\mathbf{n}_{1}(e)=\mathbf{n}(e)\mathds{1}\{e\in\pi(B)\}.$ (3.15)
Then $w(\mathbf{m})=w(\mathbf{m}_{1})w(\mathbf{m}_{2})$ and $w(\mathbf{n})=w(\mathbf{n}_{1})w(\mathbf{n}_{2})$. The current $\mathbf{n}_{2}$ is supported on $(\pi(B)\cup\pi(W))^{c}$, because $\mathbf{n}=0$ on $\pi(W)$. Also, since no edges connect to grey vertices in the exploration, the projection map $\pi$ restricts to an isomorphism between the graphs $B$ and $\pi(B)$, whose inverse we denote by $\iota:\pi(B)\to B$. Let $\tilde{\mathbf{n}}_{1}$ denote the restriction of $\hat{\mathbf{n}}_{1}$ to $B$. Then
$\iota^{-1}(\tilde{\mathbf{n}}_{1})=\mathbf{n}_{1},\quad w(\tilde{\mathbf{n}}_{1})=w(\mathbf{n}_{1}),\quad\iota^{-1}(\partial\tilde{\mathbf{n}}_{1})=\partial\mathbf{n}_{1}.$ (3.16)
Let $\mathcal{E}_{(B,W)}$ denote the event $\{(\mathfrak{B},\mathfrak{W})=(B,W)\}$, which is measurable in $\mathbf{m}_{1}+\tilde{\mathbf{n}}_{1}$ as indicated at (3.9). By conditioning on $\mathbf{m}_{2}$, $\mathbf{n}_{2}$,
$Z^{C,A}_{\mathcal{G},\mathbb{T}_{r}}[\mathcal{E}_{(B,W)}]=\sum_{\mathbf{m}_{2},\mathbf{n}_{2}}w(\mathbf{m}_{2})w(\mathbf{n}_{2})\sum_{\begin{subarray}{c}\partial\mathbf{m}_{1}=\partial\mathbf{m}_{2}\Delta C\\ \partial\mathbf{n}_{1}=\partial\mathbf{n}_{2}\Delta A\end{subarray}}w(\mathbf{m}_{1})w(\tilde{\mathbf{n}}_{1})\mathds{1}_{\mathcal{E}_{(B,W)}}\mathds{1}\{0\xleftrightarrow{(\mathbf{m}_{1}+\tilde{\mathbf{n}}_{1})_{|B}}z\},$ (3.17)
where we have $0\xleftrightarrow{(\mathbf{m}_{1}+\tilde{\mathbf{n}}_{1})_{|B}}z$ because $B\ni z$. The sum over currents $\mathbf{n}_{1}$ on $\pi(B)$ can be replaced by a sum over currents $\tilde{\mathbf{n}}_{1}$ on $B$, with the source constraint $\partial\tilde{\mathbf{n}}_{1}=\iota(\partial\mathbf{n}_{2}\Delta A)$. By Lemma 3.2 applied to the graphs $B\cup W$ and $B$,
$\sum_{\begin{subarray}{c}\partial\mathbf{m}_{1}=\partial\mathbf{m}_{2}\Delta C\\ \partial\tilde{\mathbf{n}}_{1}=\iota(\partial\mathbf{n}_{2}\Delta A)\end{subarray}}w(\mathbf{m}_{1})w(\tilde{\mathbf{n}}_{1})\mathds{1}_{\mathcal{E}_{(B,W)}}\mathds{1}\{0\xleftrightarrow{(\mathbf{m}_{1}+\tilde{\mathbf{n}}_{1})_{|B}}z\}$
$\hskip 42.67912pt=\sum_{\begin{subarray}{c}\partial\mathbf{m}_{1}=\partial\mathbf{m}_{2}\Delta C\Delta\{0,z\}\\ \partial\tilde{\mathbf{n}}_{1}=[\iota(\partial\mathbf{n}_{2}\Delta A)]\Delta\{0,z\}\end{subarray}}w(\mathbf{m}_{1})w(\tilde{\mathbf{n}}_{1})\mathds{1}_{\mathcal{E}_{(B,W)}}\mathds{1}\{0\xleftrightarrow{(\mathbf{m}_{1}+\tilde{\mathbf{n}}_{1})_{|B}}z\}.$ (3.18)
Now, we use $\partial\mathbf{n}_{1}=\iota^{-1}(\partial\tilde{\mathbf{n}}_{1})=\partial\mathbf{n}_{2}\Delta A\Delta\{0,\pi z\}$ to convert back to a sum over $\mathbf{n}_{1}$. The result is
$Z^{C,A}_{\mathcal{G},\mathbb{T}_{r}}[\mathcal{E}_{(B,W)}]=Z^{C\Delta\{0,z\},A\Delta\{0,\pi z\}}_{\mathcal{G},\mathbb{T}_{r}}[\mathcal{E}_{(B,W)}].$ (3.19)
@@
Using (3.12) and $\{x^{\prime}\in\mathcal{C}(0)\}\subset\{0\xleftrightarrow{\mathbf{m}}x^{\prime}\}$, we get
$\tau_{\beta}^{\mathbb{T}_{r}}(x)=\sum_{x^{\prime}\in\Lambda_{R}:x^{\prime}\cong x}\frac{Z^{0x^{\prime},\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[x^{\prime}\in\mathcal{C}(0)]}{Z^{\varnothing,\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}}\leq\sum_{x^{\prime}\in\Lambda_{R}:x^{\prime}\cong x}\tau_{\beta}^{\Lambda_{R}}(0,x^{\prime}).$ (3.21)
@@
The proof proceeds by bounding the terms on the right-hand side of (3.23). For each $x^{\prime}$, since $\partial\mathbf{m}=\{0,x^{\prime}\}$, there is an open path in $\mathbf{m}$ from 00 to $x^{\prime}$. When $x^{\prime}\not\in\mathcal{C}(0)$, every such path must contain a grey site. (Note that $x^{\prime}$ might be uncoloured, because the sources do not guarantee that there is an $x^{\prime\prime}\in\mathcal{C}(0)$ with $x^{\prime\prime}\cong x$.) We let $y\in{\Lambda_{R}}$ denote the last grey site on one of the paths, so that $y\xleftrightarrow{\mathbf{m}\>}x^{\prime}$ using only edges in $(\mathfrak{B}\cup\mathfrak{W})^{c}$. Also, since $y$ is grey, there exists another representative $y^{\prime}\cong y$, $y^{\prime}\neq y$ such that $y^{\prime}\in\mathcal{C}(0)$. This proves the inclusion
@@
In the following, we regard an edge subset (such as $(\mathfrak{B}\cup\mathfrak{W})^{c}$) as the graph whose edges are those in that subset and whose vertices are the vertices comprising those edges.
@@
Let $\mathcal{A}(x^{\prime},y,y^{\prime})=\{y^{\prime}\in\mathcal{C}(0)\}\cap\{y\xleftrightarrow{\mathbf{m}}x^{\prime}\text{ in }(\mathfrak{B}\cup\mathfrak{W})^{c}\}$. Then,
@@
For simplicity, we write $\mathcal{A}=\mathcal{A}(x^{\prime},y,y^{\prime})$. Since the last event in (3.25) is a connection outside of $\mathfrak{B}\cup\mathfrak{W}$, we can use it to switch sources with a third independent current without affecting $\mathfrak{B}\cup\mathfrak{W}$, as follows. Let $\bm{\ell}$ be an independent current on ${\Lambda_{R}}$ with $\partial\bm{\ell}=\varnothing$. We again condition on the event $\mathcal{E}_{(B,W)}=\{(\mathfrak{B},\mathfrak{W})=(B,W)\}$. We decompose
and similarly for $\bm{\ell}=\bm{\ell}_{1}+\bm{\ell}_{2}$. By conditioning on $\mathbf{m}_{1}$ and $\bm{\ell}_{1}$,
$Z^{\varnothing}_{\Lambda_{R}}Z^{0x^{\prime},\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{A}]=Z^{\varnothing}_{\Lambda_{R}}\sum_{B,W}Z^{0x^{\prime},\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{A}\cap\mathcal{E}_{(B,W)}]$
@@
$\quad=\sum_{B\ni y^{\prime},W}Z^{\varnothing,0x^{\prime},\varnothing}_{{\Lambda_{R}},{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{E}_{(B,W)}\cap\{y\xleftrightarrow{\mathbf{m}\>}x^{\prime}\text{ in }(B\cup W)^{c}\}]$
@@
$\quad=\sum_{B\ni y^{\prime},W}\sum_{\begin{subarray}{c}\bm{\ell}_{1},\mathbf{m}_{1}\\ \partial\mathbf{n}=\varnothing\end{subarray}}w(\bm{\ell}_{1})w(\mathbf{m}_{1})w(\mathbf{n})\mathds{1}_{\mathcal{E}_{(B,W)}}\!\!\!\!\!\!\sum_{\begin{subarray}{c}\partial\bm{\ell}_{2}=\partial\bm{\ell}_{1}\\ \partial\mathbf{m}_{2}=\partial\mathbf{m}_{1}\Delta\{0,x^{\prime}\}\end{subarray}}w(\bm{\ell}_{2})w(\mathbf{m}_{2})\mathds{1}\{y\xleftrightarrow{\mathbf{m}_{2}\>}x^{\prime}\}.$ (3.28)
@@
Since $\{y\xleftrightarrow{\mathbf{m}_{2}\>}x^{\prime}\}\subset\{y\xleftrightarrow{\bm{\ell}_{2}+\mathbf{m}_{2}\>}x^{\prime}\}$, Lemma 3.2 applied with $\mathcal{G}=\mathcal{H}=(B\cup W)^{c}$ gives
@@
$\sum_{\begin{subarray}{c}\partial\bm{\ell}_{2}=\partial\bm{\ell}_{1}\\ \partial\mathbf{m}_{2}=\partial\mathbf{m}_{1}\Delta\{0,x^{\prime}\}\end{subarray}}w(\bm{\ell}_{2})w(\mathbf{m}_{2})\mathds{1}\{y\xleftrightarrow{\mathbf{m}_{2}\>}x^{\prime}\}\leq\sum_{\begin{subarray}{c}\partial\bm{\ell}_{2}=\partial\bm{\ell}_{1}\Delta\{y,x^{\prime}\}\\ \partial\mathbf{m}_{2}=\partial\mathbf{m}_{1}\Delta\{0,y\}\end{subarray}}w(\bm{\ell}_{2})w(\mathbf{m}_{2})\mathds{1}\{y\xleftrightarrow{\bm{\ell}_{2}+\mathbf{m}_{2}\>}x^{\prime}\}.$ (3.29)
We insert this into (3.4), and use $\mathds{1}\{y\xleftrightarrow{\bm{\ell}_{2}+\mathbf{m}_{2}\>}x^{\prime}\}\leq 1$, to get
$Z^{\varnothing}_{\Lambda_{R}}Z^{0x^{\prime},\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{A}]\leq\sum_{B\ni y^{\prime},W}Z^{yx^{\prime},0y,\varnothing}_{{\Lambda_{R}},{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{E}_{(B,W)}]=Z^{yx^{\prime}}_{\Lambda_{R}}Z^{0y,\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[y^{\prime}\in\mathcal{C}(0)].$ (3.30)
The claim then follows from division by $Z^{\varnothing}_{\Lambda_{R}}$ and (3.2). ∎
@@
In the proof of Lemma 3.5, we use the notion of the backbone of a current [6, Section 4] (more on the backbone can be found in [5, 35]). A current $\mathbf{m}$ on a finite graph $\mathcal{G}$ with sources $\partial\mathbf{m}=\{u,v\}$ must have at least one path from $u$ to $v$ in the percolation configuration induced by $\mathbf{m}$. The backbone of $\mathbf{m}$ is an exploration of one of these paths. We fix some arbitrary ordering $\prec$ of the edges of $\mathcal{G}$.
@@
Let $\mathbf{m}\in\Omega_{\mathcal{G}}$ with $\partial\mathbf{m}=\{u,v\}$. The backbone of $\mathbf{m}$, denoted $\Gamma_{\mathbf{m}}$, is the unique oriented and edge self-avoiding path from $u$ to $v$, supported on edges with $\mathbf{m}$ odd, that is minimal for $\prec$. The backbone $\Gamma_{\mathbf{m}}=\{x_{i}x_{i+1}:0\leq i<k\}$ is obtained via the following exploration process:

### v2

@@ 3.2 A percolation exploration
The construction of $\mathcal{C}(0)$ is as follows. We fix an (arbitrary) ordering of the vertices in $\mathcal{G}$.
@@ 3. Step 2
3.3 A coupled cluster exploration: Proof of (2.7)
We now apply the exploration introduced in Section 3.2 to a well-chosen percolation configuration induced by random currents. This allows us to perform switching between a current on a large finite subgraph $\mathcal{G}\subset\mathbb{Z}^{d}$ and a current on $\mathbb{T}_{r}$. To proceed, we first extend the projection $\pi:\mathbb{Z}^{d}\to\mathbb{T}^{d}$ defined previously, and introduce the notion of a lift from the torus to $\mathbb{Z}^{d}$.
Projections. Given a finite or infinite subgraph $\mathcal{G}=(V,E)$ of $\mathbb{Z}^{d}$, we define its projection $\pi^{\prime}\mathcal{G}=(\pi^{\prime}V,\pi^{\prime}E)$ to be the graph on $\mathbb{T}^{d}$ with vertex set $\pi^{\prime}V=\{\pi x:x\in V\}$ and edge set $\pi^{\prime}E=\{\pi^{\prime}e:e\in E\}$, where for an edge $e=\{x,y\}\in E$ we set $\pi^{\prime}e=\{\pi x,\pi y\}$. Since the projection $\pi:\mathbb{Z}^{d}\to\mathbb{T}^{d}$ agrees with the projection $\pi^{\prime}$ applied to a graph consisting of a single vertex, we drop the prime and write simply $\pi$ for the graph projection. Thus $\pi$ can act on vertices, on edges, and on graphs. When $S$ is a subgraph of ${\mathbb{Z}^{d}}$ whose vertex set does not contain any equivalent vertices, the restriction $\pi_{|S}$ of $\pi$ to $S$ gives an isomorphism between the graphs $S$ and $\pi S$. We denote its inverse by $\pi_{|S}^{-1}$.
Lifts. Any function $F$ defined on the vertices and/or edges of the torus $\mathbb{T}^{d}$ can be lifted to the function $\hat{F}$ defined on the vertices/edges of $\mathbb{Z}^{d}$ by $\hat{F}=F\circ\pi$. In particular, given a current $\mathbf{n}$ on $\mathbb{T}^{d}$ and a subgraph $\mathcal{G}$ of $\mathbb{Z}^{d}$, the lift $\hat{\mathbf{n}}$ of $\mathbf{n}$ is the current on $\mathcal{G}$ defined by
$\hat{\mathbf{n}}=\mathbf{n}\circ\pi,\quad\text{i.e.,}\quad\hat{\mathbf{n}}(e)=\mathbf{n}(\pi e)\qquad(e\in\mathcal{G}).$ (3.11)
Let $\mathbf{m}$ be a current on $\mathcal{G}$ and $\mathbf{n}$ be a current on ${\mathbb{T}_{r}}$. The current $\mathbf{m}+\hat{\mathbf{n}}$ induces a percolation configuration $(\mathds{1}_{(\mathbf{m}+\hat{\mathbf{n}})_{e}>0})_{e\in E}$ on $\mathcal{G}$. We run the exploration of Section 3.2 on this configuration and produce $(\mathsf{B},\mathsf{G},\mathfrak{B},\mathfrak{W})$ and $\mathcal{C}(0)=(\mathsf{B},\mathfrak{B})$.
@@
Let $d\geq 1$, $r\geq 3$, $x\in\mathbb{T}_{r}$, and fix a finite graph $\mathcal{G}\supset\Lambda_{2r^{d}}$. Let $\mathbf{m}\in\Omega_{\mathcal{G}}$, and let $\mathbf{n}\in\Omega_{\mathbb{T}_{r}}$ with $\partial\mathbf{n}=\{0,x\}$. Then the exploration $(\mathsf{B},\mathsf{G},\mathfrak{B},\mathfrak{W})$ of $(\mathds{1}_{(\mathbf{m}+\hat{\mathbf{n}})_{e}>0})_{e\in E}$ satisfies the following properties:
@@
    If $e\in\mathfrak{W}$ then $\hat{\mathbf{n}}_{e^{\prime}}=0$ for every edge $e^{\prime}$ in $\mathcal{G}$ with $\pi{e^{\prime}}=\pi e$.
@@
We use the exploration to switch sources between currents on $\mathcal{G}$ and ${\mathbb{T}_{r}}$, via the following lemma.
@@
In particular, for $x\in{\mathbb{T}_{r}}$, $C=\varnothing$, $A=\{0,x\}$, and $z=x^{\prime}$ with $\pi z=x$,
@@
Let $\mathbf{m}$ denote the current on $\mathcal{G}$ and $\mathbf{n}$ denote the current on $\mathbb{T}_{r}$. Recall that $\mathcal{C}(0)$ consists of black vertices and edges. By conditioning on $(\mathfrak{B},\mathfrak{W})$, we have
@@
where $B\ni z$ means that $z$ is incident to some edge in $B$. Henceforth, we write $\mathcal{E}_{(B,W)}$ for the event $\{(\mathfrak{B},\mathfrak{W})=(B,W)\}$, which is measurable in $(\mathbf{m}+\hat{\mathbf{n}})_{|(B\cup W)}$, as indicated at (3.10). We will switch sources in $B\cup W$ (which we view as a graph).
Given an admissible pair $(B,W)$ (i.e., realisable by the exploration and such that $B\ni z$), we make the decomposition
@@
$\mathbf{n}$ $=\mathbf{n}_{1}+\mathbf{n}_{2},\qquad$ $~{}\mathbf{n}_{1}(e)=\mathbf{n}(e)\mathds{1}\{e\in\pi(B\cup W)\}.$ (3.16)
Then $w(\mathbf{m})=w(\mathbf{m}_{1})w(\mathbf{m}_{2})$ and $w(\mathbf{n})=w(\mathbf{n}_{1})w(\mathbf{n}_{2})$. We also define $\tilde{\mathbf{n}}_{1}=\hat{\mathbf{n}}_{|B\cup W}$. Since no edges connect to grey vertices in the exploration, the projection map $\pi$ restricts to an isomorphism $\pi_{|B\cup W}$ between the graphs $B\cup W$ and $\pi(B\cup W)$, and
$\tilde{\mathbf{n}}_{1}=\mathbf{n}_{1}\circ\pi_{|B\cup W},\quad w(\tilde{\mathbf{n}}_{1})=w(\mathbf{n}_{1}),\quad\pi_{|B\cup W}(\partial\tilde{\mathbf{n}}_{1})=\partial\mathbf{n}_{1}.$ (3.17)
By conditioning on $\mathbf{m}_{2}$, $\mathbf{n}_{2}$,
$Z^{C,A}_{\mathcal{G},\mathbb{T}_{r}}[\mathcal{E}_{(B,W)}\cap\{z\in B\}]$
$\qquad\qquad\qquad=\sum_{\mathbf{m}_{2},\mathbf{n}_{2}}w(\mathbf{m}_{2})w(\mathbf{n}_{2})\sum_{\begin{subarray}{c}\partial\mathbf{m}_{1}=\partial\mathbf{m}_{2}\Delta C\\ \partial\mathbf{n}_{1}=\partial\mathbf{n}_{2}\Delta A\end{subarray}}w(\mathbf{m}_{1})w(\tilde{\mathbf{n}}_{1})\mathds{1}_{\mathcal{E}_{(B,W)}}\mathds{1}\{0\xleftrightarrow{(\mathbf{m}_{1}+\tilde{\mathbf{n}}_{1})_{|B}}z\},$ (3.18)
where the connection $0\xleftrightarrow{(\mathbf{m}_{1}+\tilde{\mathbf{n}}_{1})_{|B}}z$ is present because $B\ni z$. In view of (3.17), we can change the sum over $\mathbf{n}_{1}$ to be a sum over $\tilde{\mathbf{n}}_{1}$. Now, since $\tilde{\mathbf{n}}_{1}$ and $\mathbf{m}_{1}$ are both currents on $B\cup W$, we can use Lemma 3.2 (with $\mathcal{H}=B$ and $\mathcal{E}=\mathcal{E}_{(B,W)}$) to switch sources between them. After the swiching, we use (3.17) again to change the sum over $\tilde{\mathbf{n}}_{1}$ back to a sum over $\mathbf{n}_{1}$. With $C_{1}=\partial\mathbf{m}_{2}\Delta C$ and $A_{1}=\partial\mathbf{n}_{2}\Delta A$, this process is represented by
$\sum_{\begin{subarray}{c}\partial\mathbf{m}_{1}=C_{1}\\ \partial\mathbf{n}_{1}=A_{1}\end{subarray}}=\sum_{\begin{subarray}{c}\partial\mathbf{m}_{1}=C_{1}\\ \partial\tilde{\mathbf{n}}_{1}=\pi_{|B\cup W}^{-1}A_{1}\end{subarray}}=\sum_{\begin{subarray}{c}\partial\mathbf{m}_{1}=C_{1}\Delta\{0,z\}\\ \partial\tilde{\mathbf{n}}_{1}=(\pi_{|B\cup W}^{-1}A_{1})\Delta\{0,z\}\end{subarray}}=\sum_{\begin{subarray}{c}\partial\mathbf{m}_{1}=C_{1}\Delta\{0,z\}\\ \partial\mathbf{n}_{1}=A_{1}\Delta\{0,\pi z\}\end{subarray}},$ (3.19)
where the equalities of summation symbols mean that the corresponding replacements are valid in (3.3). After reading the resulting adaptation of (3.3) from right to left, we obtain
$Z^{C,A}_{\mathcal{G},\mathbb{T}_{r}}[\mathcal{E}_{(B,W)}\cap\{z\in B\}]=Z^{C\Delta\{0,z\},A\Delta\{0,\pi z\}}_{\mathcal{G},\mathbb{T}_{r}}[\mathcal{E}_{(B,W)}\cap\{z\in B\}].$ (3.20)
@@
Using (3.13), and then dropping the event $\{x^{\prime}\in\mathcal{C}(0)\}$, we get
$\tau_{\beta}^{\mathbb{T}_{r}}(x)=\sum_{x^{\prime}\in\Lambda_{R}:x^{\prime}\cong x}\frac{Z^{0x^{\prime},\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[x^{\prime}\in\mathcal{C}(0)]}{Z^{\varnothing,\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}}\leq\sum_{x^{\prime}\in\Lambda_{R}:x^{\prime}\cong x}\frac{Z_{\Lambda_{R}}^{0x^{\prime}}Z_{\mathbb{T}_{r}}^{\varnothing}}{Z_{\Lambda_{R}}^{\varnothing}Z_{\mathbb{T}_{r}}^{\varnothing}}=\sum_{x^{\prime}\in\Lambda_{R}:x^{\prime}\cong x}\tau_{\beta}^{\Lambda_{R}}(0,x^{\prime}).$ (3.22)
@@
The proof proceeds by bounding the terms on the right-hand side of (3.24). These terms involve a current $\mathbf{m}$ on $\Lambda_{R}$ with $\partial\mathbf{m}=\{0,x^{\prime}\}$, so there is an open path in $\mathbf{m}$ from 00 to $x^{\prime}$. When $x^{\prime}\not\in\mathcal{C}(0)$, every such path must contain a grey site. (Note that $x^{\prime}$ might be uncoloured, because the sources do not guarantee that there is an $x^{\prime\prime}\in\mathcal{C}(0)$ with $x^{\prime\prime}\cong x$.) We let $y\in{\Lambda_{R}}$ denote the last grey site on one of the paths from 00 to $x^{\prime}$, so that $y\xleftrightarrow{\mathbf{m}\>}x^{\prime}$ using only edges in $(\mathfrak{B}\cup\mathfrak{W})^{c}$. Also, since $y$ is grey, there exists another representative $y^{\prime}\cong y$, $y^{\prime}\neq y$ such that $y^{\prime}\in\mathcal{C}(0)$. This proves the inclusion
@@
As in the proof of Lemma 3.4, we regard an edge subset (such as $(\mathfrak{B}\cup\mathfrak{W})^{c}$) as the graph whose edges are those in that subset and whose vertices are the vertices comprising those edges.
@@
Let $\mathcal{A}(x^{\prime},y,y^{\prime})=\{y^{\prime}\in\mathcal{C}(0)\}\cap\{y\xleftrightarrow{\mathbf{m}\>}x^{\prime}\text{ in }(\mathfrak{B}\cup\mathfrak{W})^{c}\}$ denote the event on the right-hand side of (3.26). Then,
@@
The left-hand side of the desired inequality involves two currents, but we will need a third for the two-point function on the right-hand side. For simplicity, we write $\mathcal{A}=\mathcal{A}(x^{\prime},y,y^{\prime})$. Since the connection joining $y$ and $x^{\prime}$ is a connection outside of $\mathfrak{B}\cup\mathfrak{W}$, we can use it to switch sources with a third independent current without affecting $\mathfrak{B}\cup\mathfrak{W}$, as follows. We again condition on the event $\mathcal{E}_{(B,W)}=\{(\mathfrak{B},\mathfrak{W})=(B,W)\}$ and make the decomposition
$\mathbf{m}=\mathbf{m}_{1}+\mathbf{m}_{2},\qquad\mathbf{m}_{1}(e)=\mathbf{m}(e)\mathds{1}\{e\in B\cup W\}.$ (3.28)
Let $\bm{\ell}$ be an independent current on $(B\cup W)^{c}$ with $\partial\bm{\ell}=\varnothing$. By conditioning on $\mathbf{m}_{1}$, we have
$Z^{0x^{\prime},\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{A}]=\sum_{B,W}Z^{0x^{\prime},\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{A}\cap\mathcal{E}_{(B,W)}]$
@@
$\quad=\sum_{B\ni y^{\prime},W}Z^{0x^{\prime},\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{E}_{(B,W)}\cap\{y\xleftrightarrow{\mathbf{m}\>}x^{\prime}\text{ in }(B\cup W)^{c}\}]\frac{Z^{\varnothing}_{(B\cup W)^{c}}}{Z^{\varnothing}_{(B\cup W)^{c}}}$
@@
$\quad=\sum_{B\ni y^{\prime},W}\sum_{\begin{subarray}{c}\mathbf{m}_{1}\\ \partial\mathbf{n}=\varnothing\end{subarray}}w(\mathbf{m}_{1})w(\mathbf{n})\mathds{1}_{\mathcal{E}_{(B,W)}}\!\!\!\!\!\!\sum_{\begin{subarray}{c}\partial\bm{\ell}=\varnothing\\ \partial\mathbf{m}_{2}=\partial\mathbf{m}_{1}\Delta\{0,x^{\prime}\}\end{subarray}}w(\bm{\ell})w(\mathbf{m}_{2})\mathds{1}\{y\xleftrightarrow{\mathbf{m}_{2}\>}x^{\prime}\}\frac{1}{Z^{\varnothing}_{(B\cup W)^{c}}}.$ (3.29)
@@
We then switch sources between $\bm{\ell}$ and $\mathbf{m}_{2}$ on $\mathcal{G}=\mathcal{H}=(B\cup W)^{c}$. Since $\{y\xleftrightarrow{\mathbf{m}_{2}\>}x^{\prime}\}\subset\{y\xleftrightarrow{\bm{\ell}+\mathbf{m}_{2}\>}x^{\prime}\}$, from (3.7) we get
@@
$\sum_{\begin{subarray}{c}\partial\bm{\ell}=\varnothing\\ \partial\mathbf{m}_{2}=\partial\mathbf{m}_{1}\Delta\{0,x^{\prime}\}\end{subarray}}w(\bm{\ell})w(\mathbf{m}_{2})\mathds{1}\{y\xleftrightarrow{\mathbf{m}_{2}\>}x^{\prime}\}\leq\sum_{\partial\mathbf{m}_{2}=\partial\mathbf{m}_{1}\Delta\{0,y\}}w(\mathbf{m}_{2})\sum_{\partial\bm{\ell}=\{y,x^{\prime}\}}w(\bm{\ell}).$ (3.30)
After insertion of (3.30) into (3.4), the sum over $\bm{\ell}$, together with $1/Z^{\varnothing}_{(B\cup W)^{c}}$, creates a two-point function on $(B\cup W)^{c}$ which is bounded by $\tau_{\beta}^{\Lambda_{R}}(y,x^{\prime})$ by a Griffiths inequality. The remaining factors reassemble to provide the bound
$Z^{0x^{\prime},\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{A}]\leq\tau_{\beta}^{\Lambda_{R}}(y,x^{\prime})\sum_{B\ni y^{\prime},W}Z^{0y,\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[\mathcal{E}_{(B,W)}]=\tau_{\beta}^{\Lambda_{R}}(y,x^{\prime})Z^{0y,\varnothing}_{{\Lambda_{R}},{\mathbb{T}_{r}}}[y^{\prime}\in\mathcal{C}(0)],$ (3.31)
as desired. ∎
@@
In the proof of Lemma 3.5, we use the notion of the backbone of a current [6, Section 4] (more on the backbone can be found in [5, 37, 38]). A current $\mathbf{m}$ on a finite graph $\mathcal{G}$ with sources $\partial\mathbf{m}=\{u,v\}$ must have at least one path from $u$ to $v$ in the percolation configuration induced by $\mathbf{m}$. The backbone of $\mathbf{m}$ is an exploration of one of these paths. We fix some arbitrary ordering $\prec$ of the edges of $\mathcal{G}$.
@@
Let $\mathbf{m}\in\Omega_{\mathcal{G}}$ with $\partial\mathbf{m}=\{u,v\}$. The backbone of $\mathbf{m}$, denoted $\Gamma_{\mathbf{m}}$, is the unique oriented and edge self-avoiding path from $u$ to $v$, supported on edges $e$ with $\mathbf{m}_{e}$ odd, that is minimal for $\prec$. The backbone $\Gamma_{\mathbf{m}}=\{x_{i}x_{i+1}:0\leq i<k\}$ is obtained via the following exploration process: